OpenStudy (anonymous):

Find all the points of inflection of the function. y=x^1/3(x-4)

myininaya (myininaya):

So your goal is to find the second derivative. I will give you a new goal when you have done that.

OpenStudy (anonymous):

the second derivative is what i feel I keep getting wrong because for the first I get \[x^{1/3}+\frac{ 1 }{ 3 } x^2 -4\] and for the second i get \[\frac{ 2 }{ 3 }x ^{4/3}+\frac{ 1 }{ 9 }x^2-\frac{ 4 }{ 3 }\]

OpenStudy (anonymous):

I feel like thats wrong

myininaya (myininaya):

Is your y this: \[y=x^\frac{1}{3}(x-4)?\]

OpenStudy (anonymous):

no its \[x ^{1/3}(x-4)\]...the 1/3 is the exponent

myininaya (myininaya):

Ok so we used the product rule: \[y'=x^\frac{1}{3} \cdot (1-0)+\frac{1}{3}x^{\frac{1}{3}-1} \cdot (x-4)\] Is this what you did for the first derivative?

myininaya (myininaya):

\[y'=x^\frac{1}{3}(1)+\frac{1}{3}x^\frac{-2}{3}(x-4)\] \[y'=x^\frac{1}{3}+\frac{1}{3}x^\frac{1}{3}-\frac{4}{3}x^\frac{-2}{3}\] ?

OpenStudy (anonymous):

yea thats what I got the first time I tried the problem

myininaya (myininaya):

I don't you did. Unless you typed the wrong thing.

myininaya (myininaya):

Anyways we do have like terms in y'. Combine those.

OpenStudy (anonymous):

no, what I'm trying to say is that the answer you posted earlier was the same as what I got when I redid the first derivative

OpenStudy (anonymous):

so Now i just need to find the second derivative

myininaya (myininaya):

yep so you are trying to find the derivative of this: \[y'=\frac{4}{3}x^\frac{1}{3}-\frac{4}{3}x^\frac{-2}{3}\]

OpenStudy (anonymous):

ok so then I get \[\frac{ 4 }{ 9 }x ^{-2/3}+ \frac{ 8 }{ 9 }x^{-5/3}\].....right

myininaya (myininaya):

right

myininaya (myininaya):

Objective write as one fraction Main objective to find when y'' is 0 and y'' dne. \[y''=\frac{4}{9x^\frac{2}{3}}+\frac{8}{9x^\frac{5}{3}}\] Combine fractions: \[y''=\frac{4 x^\frac{3}{3}}{9x^\frac{2}{3} x^\frac{2}{3}}+\frac{8}{9x^\frac{5}{3}}\]

myininaya (myininaya):

\[y''=\frac{4x+8}{9x^\frac{5}{3}}\]

OpenStudy (anonymous):

ok so now i set it equal to zero right

OpenStudy (anonymous):

thanks

myininaya (myininaya):

Main objective to find when y'' is 0 and y'' dne. Then you see if the concavity switchs at both of those numbers.